OpenStudy (anonymous):

how to prove (a_n+b_n)/2 > sqrt(a_n*b_n) using induction?

OpenStudy (turingtest):

I'm not the master of induction, but I can try... are there any restrictions on \(a_n\) and \(b_n\) ?

OpenStudy (turingtest):

@eliassaab that's a nice generalization, but not a proof by induction

OpenStudy (anonymous):

a_n+1=1/2(a_n +b_n) b_n+1=sqrt(a_n x b_n)

OpenStudy (anonymous):

i understood the relation between those two. but i cant solve the proof using induction

OpenStudy (anonymous):

Read this from wikipeia In two dimensions, 2x1 + 2x2 is the perimeter of a rectangle with sides of length x1 and x2. Similarly, 4√(x1x2) is the perimeter of a square with the same area. Thus for n = 2 the AM-GM inequality states that a square has the smallest perimeter among all rectangles with equal area,

OpenStudy (turingtest):

again, not inductive...

OpenStudy (anonymous):

@eliassaab , tq for that. i understand about that am-gm theory, but i need the induction part of proving

OpenStudy (anonymous):

For any two positive numbers a and b \[ \sqrt{a b} \le \frac { a+b} 2 \] There is no need for induction here.

OpenStudy (anonymous):

ok i see that this is probably my fault, because i mentioned induction earlier. forget that. it is always the case that \[\frac{a+b}{2}\geq \sqrt{ab}\] the subscript n is irrelevant

OpenStudy (turingtest):

but that's what they asked for...

OpenStudy (turingtest):

is this a follow up Q ?

OpenStudy (anonymous):

no i think that was because i stupidly mentioned it in the answer to a previous problem, which his what i get for writing before thinking clearly

OpenStudy (turingtest):

ah, ic

OpenStudy (anonymous):

forget i mentioned it, it was stupid on my part

OpenStudy (anonymous):

nobody is stupid. we learn n learn....

OpenStudy (anonymous):

@TuringTest original question was here http://openstudy.com/study#/updates/4f8eb037e4b000310face9e5

OpenStudy (anonymous):

means, we can just explain like that, and no need for induction part?